# Knowledge-guided machine learning for county-level corn yield prediction under drought

### v2

@@ Abstract
Remote sensing (RS) technique, enabling the non-contact acquisition of extensive ground observations, is a valuable tool for crop yield predictions. Traditional process-based models struggle to incorporate large volumes of RS data, and most users lack understanding of crop growth mechanisms. In contrast, machine learning (ML) models are often criticized as “black boxes” due to their limited interpretability. To address these limitations, we utilized Knowledge-Guided Machine Learning (KGML), a framework that leverages the strengths of both process-based and ML models. Existing works have either overlooked the role of soil moisture in corn growth or did not embed this effect into their models. To bridge this gap, we developed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, treating soil moisture as an intermediate variable in corn growth to emphasize its key role in plant development. Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalized predicted yield in drought-affected areas. Our experiments showed that the KGML-SM model outperformed other traditional ML models. We explored the relationships between drought, soil moisture, and corn yield prediction by assessing the importance of different features within the model, and analyzing how soil moisture impacts predictions across different regions and time periods. Finally we provided interpretability for prediction errors to guide future model optimization.
@@
Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought-aware learning
journal: Journal
@@
[inst3]organization=Research and Development Division, National Agricultural Statistics Service, United States Department of Agriculture,addressline=Washington, postcode=20250, state=DC, country=USA
@@ 1 Introduction
Corn, as a primary crop, plays a vital role in the U.S. agriculture, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969). Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005). Achieving accurate prediction across large areas and diverse environmental conditions is challenging (Weiss et al., 2020). Traditional process-based models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004). These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021). Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and the Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010). Many studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields. However, process-based models have several limitations. Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process. Moreover, process-based models require extensive manual parameter tuning, limiting their applicability for applications across vast regions.
In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015). It enables the collection of large-scale real-time data across vast agricultural areas. This data, typically combined with machine learning (ML), facilitates accurate and efficient large-scale yield prediction. Deep learning (DL) (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging a large amount of RS data by automatically learning features from big datasets. This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between crop environmental factors and crop yield (He et al., 2016). Many studies applying DL to crop yield prediction have shown its advantages in performance improvement. You et al. (2017) introduced a scalable approach to crop yield prediction using modern representation learning rather than traditional hand-crafted features, achieving impressive results. Wang et al. (2018) demonstrated soybean yield predictions in Argentina and obtained reliable results in Brazil using transfer learning, even with limited data. Ma et al. (2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks. However, despite the popularity of DL models in corn yield prediction tasks, they have certain limitations. First, DL models require a large amount of data to learn the relationships between inputs and outputs. Additionally, DL models often function as a black box, lacking reliable interpretability of the prediction results. Specifically, crop growth involves complex biological processes among soil, weather, and corn, which are challenging for DL models to fully capture. Therefore, how to leverage advantages of process-based models and DL models to address these challenges has become an important research topic.
Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better performance, scientific consistency, and explainability of results. KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining. Knowledge-Guided Learning incorporates scientific knowledge, such as physical laws, into the learning algorithm by modifying the loss function. It guides the optimization process toward scientifically consistent and generalizable solutions (Daw et al., 2022; Bao et al., 2021). Knowledge-Guided Architecture embeds scientific knowledge directly into the structure of the ML model. By designing knowledge-guided neural network architectures, researchers can enable models to automatically capture physical properties or conservation laws, which improves both interpretability and robustness (Dugdale et al., 2017; Luo et al., 2023). Knowledge-Guided Pretraining uses simulated data or self-supervised learning tasks to pretrain ML models. By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023). Building on these approaches, KGML has been increasingly used in various domains, such as agriculture, hydrology, and biology, enabling improved data-driven decision-making and predictive modeling.
Recently, many studies have used KGML to combine process-based and ML models for crop yield predictions (Kimball et al., 2023; Burroughs et al., 2023). Since process-based models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023). These studies aim to identify patterns between agricultural features and crop yields from large datasets using ML, while leveraging data simulated by process-based models to provide interpretability. Specifically, intermediate variables in crop growth generated by process-based models are used to train ML models, allowing the ML models to simulate specific parts of the crop growth process. For example, He et al. (2023) integrated physical knowledge from process-based models and enabled the model to adjust for temporal data shifts over the years. Yang et al. (2023) used both historical and in-season RS data to better predict key agricultural variables such as grain yield and carbon cycling processes. Yang et al. (2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield predictions have overlooked the critical role of soil, particularly how soil moisture contributes to corn growth and affects yield. Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth. As a key indicator of crop water supply, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998). Unlike most RS data, such as NDVI, which only reflects the growth status of plants at a specific moment, soil moisture has a cumulative impact on yield over time. Nonetheless, a significant challenge for ML models is the limited availability of soil moisture data. For example, the widely used SPL4SMGP.007 SMAP soil moisture dataset (Reichle et al., 2022) only covers the period from 2015 onward, which limits the ability of ML models to learn from extensive historical data and impacts model performance. Therefore, we present a Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework that explicitly incorporated the soil moisture impact in the ML model and provided interpretability. The model consisted of two major components: a Weather-to-Soil (W2S) encoder modeling the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) integrating features from multiple sources and dynamically weighting their influences on corn yield across time and space. To train this model, we adopted a two-step process. First, we used the APSIM model to generate a simulated field-level dataset for pretraining. Then, the model was finetuned using a county-level dataset from Google Earth Engine (GEE) (Gorelick et al., 2017) and United States Department of Agriculture (USDA) National Agricultural Statistics Service (NASS).
Furthermore, we conducted research into the relationship between soil moisture and corn yield prediction in drought-affected areas, where drought often reduced yields and caused overestimation (Cooper et al., 2014). Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017). However, these studies either had a narrow scope, or did not utilize this connection for model optimization. we designed a drought-aware loss function to further enhance model performance under drought conditions and mitigate overestimation problem. Finally, based on the relationship between soil moisture and corn yield prediction, we provided explanations for the prediction errors of the KGML-SM model and offered directions for future model optimization. Our research aimed to improve the accuracy and interpretability of corn yield predictions, ultimately contributing to more resilient agricultural systems under extreme weather variability.
Our main contributions are as follows:
  • 1.
    This paper introduced the KGML-SM framework. This framework integrated process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
  • 2.
    A drought-aware loss function was designed. This function enhanced model performance under drought conditions and mitigated overestimation.
  • 3.
@@ 2 Data acquisition
In this study, we developed two datasets for corn yield prediction. The first dataset was a field-level dataset generated using APSIM (McCown et al., 1996) and used for pretraining. The second dataset was a county-level dataset derived from GEE (Gorelick et al., 2017) and USDA NASS and used for finetuning. The workflow was divided into three main steps: simulation, pretraining, and finetuning. A summary of all public datasets used in this study is provided in Table 1. In this section, we first introduce the study area (Sec. 2.1); then, we provide details of the APSIM field-level dataset (Sec. 2.2); finally, we describe the construction of the GEE county-level dataset (Sec. 2.3).
Table 1: Public datasets used in this study. a, b, and c indicate different purposes: a for APSIM simulation, b for model pretraining, and c for model finetuning.
Dataset Use Description References
Iowa Environmental Mesonet a, b In-situ weather data (Herzmann et al., 2004)
PRISM c Gridded weather data (Daly et al., 2015, 2008)
MODIS MCD43A4 c Vegetation indices (Schaaf and Wang, 2015)
MODIS MCD18A1.061 c Radiation data (Wang, 2021)
USDA-CDL c Crop data layer (USDA-NASS, 2017)
SMAP c Soil moisture data (Reichle et al., 2022)
USDA NASS yield c County-level yield (USDA, 2020)
@@ 2.1 Study area
Our research focused on corn yield prediction across the U.S. Corn Belt. Twelve states are selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. These states are crucial agricultural states in the U.S., known for their significant contributions to corn production. We generated a five-year average yield map (Fig. 1) for these twelve states and considered them well suited for corn yield prediction research.

### v3

@@ Abstract
Remote sensing (RS) provides large-scale, non-contact observations that are valuable for crop yield prediction. Process-based models are based on crop growth mechanisms but often struggle with large RS datasets and require heavy calibration. Machine learning (ML) models can use RS data effectively but are often criticized as “black boxes” with low interpretability. To overcome these issues, we proposed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, which combined the strengths of both approaches. Soil moisture was used as an intermediate variable, linking weather and crop growth and serving as a key factor in drought impacts on yield. This design improved interpretability by tracing yield prediction errors to soil moisture estimates. We also added a drought-aware loss function that penalized overestimation in dry regions, making the model more robust under drought stress. This study focused on the U.S. Corn Belt, covering 12 states and more than 800 counties from 2019 to 2023. We constructed two datasets: a field-level dataset generated from the Agricultural Production Systems sIMulator (APSIM) and a county-level dataset built with satellite-based MODIS RS data and gridded climate data. Model performance was evaluated against USDA-NASS county-level yield records. KGML-SM achieved lower errors than traditional ML baselines, with an RMSE of 1.071 t/ha and an $R^{2}$ of 0.807 in 2023. Attention-based analysis further revealed the role of drought and soil moisture in yield prediction. Overall, KGML-SM improves both accuracy and interpretability, offering insights for future model development and climate-resilient agriculture.
@@
Crop modeling; Remote sensing; Process-based models; Attention mechanism; Drought-aware loss function; Agricultural informatics
journal: Remote Sensing of Environment
@@
[inst3]organization=U.S. Department of Agriculture, National Agricultural Statistics Service, city=Washington, postcode=20250, state=DC, country=USA
[…]
[inst4]organization=U.S. Department of Agriculture, Agricultural Research Service, Genetics and Sustainable Agricultural Research Unit, city=Mississippi State, postcode=39762, state=MS, country=USA
[…]
[inst5]organization=NASA Goddard Space Flight Center, city=Greenbelt, postcode=20771, state=MD, country=USA
@@ 1 Introduction
Corn, as a primary crop, plays a vital role in U.S. agriculture, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969). Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005). Nevertheless, achieving accurate prediction across large areas and diverse environmental conditions remains challenging (Lobell et al., 2015). Climate extremes such as drought pose a major obstacle to accuracy, particularly at the county level where heterogeneous management practices and soil conditions introduce additional variability (Rigden et al., 2020).
@@
Traditional process-based models are grounded in a detailed understanding of the physical, chemical, and biological processes of crop growth (Puntel et al., 2016). These models use equations derived from scientific principles to simulate crop, weather, and soil interactions, which makes them highly interpretable and reliable when underlying processes are well characterized (Shahhosseini et al., 2021). Prominent frameworks include the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the Agricultural Policy/Environmental eXtender (APEX) (Williams and Izaurralde, 2010), and the Ecosys Modelling Project (ecosys) (Grant, 2001), all of which have been successfully applied to crop yield prediction (Zhen et al., 2023, 2022; Asseng et al., 2013; Schauberger et al., 2017). Yet plant growth is highly complex, and the reliance on a limited set of fixed inputs prevents these models from fully capturing this complexity (Lobell et al., 2015; Müller et al., 2017). In addition, they require extensive manual parameter tuning, which constrains their use across large regions (Kimball et al., 2023; Zhen et al., 2022).
In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015). It enables the collection of large-scale real-time data across vast agricultural areas. These data, typically combined with machine learning (ML), facilitate accurate and efficient large-scale yield prediction. Deep learning (DL), a subset of ML, is based on neural networks and is particularly effective for high-dimensional RS data, as it can automatically extract features from large datasets and capture complex environment–yield relationships without manual engineering (Goodfellow et al., 2016; He et al., 2016). Several studies have demonstrated the potential of DL for yield prediction, such as scalable representation learning (You et al., 2017), transfer learning across regions (Wang et al., 2018), and county-level prediction using Bayesian neural networks (Ma et al., 2021a). More recently, adaptive multi-modal fusion frameworks have been proposed to integrate heterogeneous data sources and provide improved accuracy (Mena et al., 2025). However, DL methods face important limitations: they require large amounts of training data, often function as black boxes with limited interpretability, and struggle to represent the biological processes linking soil, weather, and crop growth. In contrast, process-based models explicitly capture these mechanisms but are less suited for large-scale RS integration. This complementarity highlights the need for approaches that combine the strengths of both, motivating recent research on combining these two kinds of models. A recent study (Sadeh et al., 2024) also emphasized that relying solely on either process-based or ML models is insufficient for robust yield prediction. This growing recognition of complementarity has motivated the emergence of knowledge-guided machine learning (KGML), which seeks to integrate process-based understanding into ML frameworks to enhance both accuracy and interpretability.
KGML (Karpatne et al., 2022) aims to integrate scientific knowledge into ML frameworks to achieve better performance, scientific consistency, and explainability of results. This paradigm involves three main approaches: (i) knowledge-guided learning, which incorporates scientific laws into algorithms through modified loss functions (Daw et al., 2022; Bao et al., 2021); (ii) knowledge-guided architectures, which embed knowledge directly into model structures (Dugdale et al., 2017; Luo et al., 2023); and (iii) knowledge-guided pretraining, which uses simulated data or self-supervised tasks to initialize models (Licheng et al., 2022; Chen et al., 2023). In crop yield prediction, most KGML applications rely on DL rather than traditional ML approaches (Burroughs et al., 2023; He et al., 2023; Yang et al., 2023). Since process-based models already provide insights into crop growth and generate high-quality simulated data, Knowledge-Guided Pretraining has become the most popular strategy. These studies typically use ML to learn yield-relevant patterns while incorporating simulated intermediates as mechanistic links from weather and soil to crop growth (He et al., 2023; Yang et al., 2023, 2024). Specifically, He et al. (2023) applied physics-guided reweighting with ecosystem-process intermediates—ecosystem autotrophic respiration (Ra), ecosystem heterotrophic respiration (Rh), and net ecosystem exchange (NEE)—to improve time-aware robustness and interpretability. Yang et al. (2023) coupled process-model surrogates with Ensemble Kalman Filter (EnKF) data assimilation to fuse leaf area index (LAI), gross primary production (GPP), and evapotranspiration (ET) for yield prediction. Yang et al. (2024) linked image-derived plant traits to a knowledge-guided S-shaped growth curve to forecast fruit growth, yield, and maturity. Recently, Miranda et al. (2024) developed a physics-informed recurrent neural network for yield loss forecasting, where crop water use (ETa) and drought sensitivity (Ky) were estimated under physical constraints.
Despite these advances, no study has explicitly embedded soil moisture into the ML model structure. Soil moisture is central to crop production because it mediates the effects of weather on plant growth and yield formation (Unganai and Kogan, 1998). Unlike vegetation indices (VIs) which capture plant status at a single moment, soil moisture reflects the cumulative effects of precipitation (PPT) and temperature over time, making it a more process-oriented indicator of water availability. Embedding soil moisture explicitly in a KGML model not only provides a mechanistic link between weather and yield but also creates diagnostic interpretability, allowing prediction errors to be traced to potential misrepresentation of soil water dynamics.
To effectively exploit soil moisture for yield prediction, reliable large-scale observations are needed. Several soil moisture products have been developed over the past decades and provide valuable long-term global records. The ASCAT product offers C-band scatterometer-based soil moisture observations (Wagner et al., 2013). In addition to satellite scatterometer data, reanalysis products such as ERA5-Land (Muñoz-Sabater et al., 2021) and MERRA-2 (Gelaro et al., 2017) provide surface and root-zone soil moisture at global scales. Furthermore, the ESA product delivers more than 30 years of consistent global daily soil moisture records by merging multiple satellite sensors (Dorigo et al., 2017). Building on these advances, the Soil Moisture Active Passive (SMAP) mission (Reichle et al., 2022; Entekhabi et al., 2010) delivers high-accuracy global soil moisture observations with a revisit cycle of two to three days. Compared with earlier C- or X-band missions, SMAP’s L-band measurements penetrate deeper into the soil and are less affected by vegetation and atmospheric disturbances, making it one of the most reliable datasets for agricultural applications.
Previous studies consistently demonstrate a strong link between soil moisture, corn yield, and drought impacts. For instance, Ines et al. (2013a) showed that assimilating remotely sensed soil moisture and VIs into a crop simulation model substantially improved corn yield prediction. In addition, Rigden et al. (2020) reported that the combined influence of soil moisture and atmospheric evaporative demand explained most of the observed interannual variability in U.S. corn yields, underscoring the need to jointly consider soil and atmospheric drivers. At the field scale, de Lara et al. (2019) found that soil water content in deeper layers during reproductive stages showed the strongest correlation with corn yield, and that integrating high-resolution VIs improved precision irrigation decisions. Other studies have further explored related directions, such as improving retrieval accuracy by combining ASCAT and SMAP observations (Ma et al., 2024), developing cumulative drought indices (CDI) from process-based models to enhance subfield yield predictions (Shuai and Basso, 2022), or applying ML and DL approaches for soil moisture prediction and yield modeling (Teshome et al., 2024; Kang et al., 2020; Yu et al., 2020). Collectively, these works (Table 1) underscore the central role of soil moisture in corn production, but most treat it only as an auxiliary covariate or prediction target. In contrast, our KGML-SM framework embeds soil moisture explicitly as an intermediate variable, creating a mechanistic link between weather and yield and enabling diagnostic interpretability. Building on this design, KGML-SM also introduces a drought-aware loss function to penalize overestimation under water-limited conditions, thereby improving robustness. Unlike approaches that simply add soil moisture as another input feature or prediction, our framework provides a structured way to trace yield errors back to soil–weather interactions. Moreover, the proposed modules are general: embedding soil moisture as an intermediate variable can be integrated into other crop yield models to enhance interpretability, while the drought-aware loss can be applied more broadly in drought-prone regions.
Table 1: Summary of KGML applications in crops and soil moisture–related studies in context of this work.
Reference Content Method
(He et al., 2023) Using physics-guided neural networks for time-aware fairness in crop yield prediction LSTM with attention; physics-guided reweighting; fairness refinement
(Yang et al., 2023) Integrating process-based surrogates with multi-source data assimilation for agroecosystem prediction GRU surrogate; EnKF; data fusion
(Yang et al., 2024) Predicting greenhouse strawberry growth trajectory and yield via knowledge-guided computer vision Faster R-CNN; DenseNet-based trait extraction; S-shaped growth-curve modeling
(Lobell et al., 2015) Mapping large-area crop yield from satellites via pseudo-observations and calibration Pseudo-observation conversion; regression calibration
(Shuai and Basso, 2022) Improving subfield corn yield prediction by including in-season water deficit with RS VIs SALUS-simulated CDI by yield-stability zones; Random Forest (RF) per tile-year and a two-window composite
(Mena et al., 2025) Fusing multi-modal remote-sensing data for optimal subfield yield prediction Multi-modal encoders; gated fusion
(Miranda et al., 2024) Exploring physics-informed neural networks for crop yield loss forecasting PINN with physics-based constraints
(Ines et al., 2013a) Assimilating RS soil moisture and vegetation into a crop model for corn yield prediction EnKF with DSSAT-CSM (Maize); sequential SM/LAI assimilation
(Ma et al., 2024) Integrating ASCAT and SMAP for global surface soil moisture retrieval using ML Model comparison: RF / LSTM / SVM / CNN
(Rigden et al., 2020) Combining soil moisture and atmospheric evaporative demand to predict US corn yields Statistical modeling with process-based analysis
(de Lara et al., 2019) Relating soil water content and high-resolution imagery for precision irrigation in corn Linear regression with VIs coupling
(Teshome et al., 2024) Benchmarking DL vs. ML models for soil moisture prediction under irrigation treatments
(Kang et al., 2020) Comparing environmental variables and ML algorithms for corn yield in the US Midwest Comparative ML assessment
(Yu et al., 2020) Predicting multi-depth soil water content during summer corn for irrigation planning Residual bidirectional LSTM
This study The first study to embed soil moisture explicitly as an intermediate variable within the ML architecture for county-level corn yield prediction; mitigating overestimation under drought area Knowledge-guided pretraining; W2S encoder; drought-aware loss function
In summary, this study introduces a KGML-SM framework to address the limitations of existing process-based and ML approaches for county-level corn yield prediction. The framework explicitly embeds soil moisture as an intermediate variable through a Weather-to-Soil (W2S) encoder and an attention mechanism (Vaswani, 2017), creating a mechanistic link between weather drivers and crop outcomes. In addition, a drought-aware loss function penalizes overestimation under water-limited conditions, thereby improving model robustness during drought years. By jointly enhancing accuracy, interpretability, and robustness, KGML-SM improves yield prediction. To our knowledge, this is the first framework to explicitly embed soil moisture into a KGML model for county-level prediction, offering both methodological innovation and practical relevance for supporting resilient agricultural management under increasing climate variability.
@@ 2 Data acquisition
In this study, we developed two datasets for corn yield prediction. The first dataset was a field-level dataset generated using APSIM (McCown et al., 1996) and used for pretraining. The second dataset was a county-level dataset derived from Google Earth Engine (GEE) (Gorelick et al., 2017) and USDA NASS (USDA, 2020), and used for finetuning. The workflow was divided into three main steps: simulation, pretraining, and finetuning. In this section, we first introduce the study area (Sec. 2.1); then, we provide details of the APSIM field-level dataset (Sec. 2.2); finally, we describe the construction of the GEE county-level dataset (Sec. 2.3).
@@ 2.1 Study area
Our research focused on corn yield prediction across the U.S. Corn Belt. Twelve states were selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. These states are crucial agricultural states in the U.S., known for their significant contributions to corn production. We generated a five-year average yield map (Fig. 1) for these twelve states and considered them well suited for corn yield prediction research. For brevity, state names were referred to by their standard abbreviations (e.g., Wisconsin as WI) in the following sections.